OpenStudy (anonymous):

evaluate the logarithmic equation ln e ^ 2

OpenStudy (lgbasallote):

use power rule..you know how to do that right?

OpenStudy (anonymous):

\[\log_b(b^x)=x\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and \[\log_e(z)=\ln(z)\]

OpenStudy (lgbasallote):

power rule still works @satellite73 :P

OpenStudy (lgbasallote):

and \(\ln e = 1\)

OpenStudy (anonymous):

you are composing a function with its inverse \[f(x)=e^x,f^{-1}(x)=\ln(x)\] so \[\ln(e^x)=x\] and also \[e^{\ln(x)}=x\]

OpenStudy (anonymous):

isn't it 2e^2

OpenStudy (anonymous):

no, it is 2

OpenStudy (anonymous):

i put that the answer was 1 on my test and i got it wrong...

OpenStudy (lgbasallote):

\(\ln e^{x} = x\ln e = x\) my way :P haha

OpenStudy (anonymous):

but the exponent doesn't change...according to the rules.....at least that's what I remember....but maybe I'm wrong....

OpenStudy (anonymous):

\(\ln(e^x)=x\) and so \(\ln(e^2)=2\) just like \(\log_5(5^2)=2\)

OpenStudy (anonymous):

could the answer just be e?

OpenStudy (anonymous):

so its 2?

OpenStudy (anonymous):

yes it is 2

OpenStudy (lgbasallote):

yes 2

OpenStudy (anonymous):

ok so how did you find that... cause im stupid

OpenStudy (anonymous):

you are asking the following question: what number would you raise e to in order to get \(e^2\) and the answer is clearly 2

OpenStudy (anonymous):

as i wrote above, it is always the case that \(\log_b(b^x)=x\)

OpenStudy (lgbasallote):

there are 2 ways... satellite's way is the logarithmic propert \(\ln e^x = x\) my way is \(\ln e^x = x \ln e = x\)

OpenStudy (anonymous):

it is a tautology, like asking who is buried in grants tomb

OpenStudy (anonymous):

\(\log_{10}(10^6)=6\) etc